OpenStudy (anonymous):

how do i graph using the x and y intercepts.....6y=4x-18

OpenStudy (aroub):

x-inter =(9/2,0) y-inter=(0,-3)

OpenStudy (aroub):

if u want to find the x-intercept u put y=0 and when u want to find the y-intercept u put x=0

OpenStudy (anonymous):

I have a hard time trying to figuire the steps out, do you have an easier way

OpenStudy (aroub):

this is the easiest way i know , just when u want to find the x-intercept u put y=0 and when u want to find y-intercept u put y=0 <-- did u understand??